Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Nigh report 2013-12-10

Observers: Michel Dennefeld, Joe Lyman, David Bersier        TO: Francisco
Support team: M. Fraser, S. Taubenberger  
TAT on duty:  Joe Anderson, Morgan Fraser
------------------------------------------------

Instruments: EFOSC2

WEATHER CONDITIONS
beginning: wind speed 7 m/s; humidity 42 % ; clear
middle: wind speed 3 m/s; humidity 30 %;  clear
end: wind speed 2 m/s; humidity 25 %;  clear

Afternoon CALIBRATIONS: EFOSC bias, dome flats, arcs

Sky flats: UVi

OBSERVATIONS:  official night start at UT01:10 (=LST1:47)

Target                              Type             UT         Filt/Gm     Slit      Grade    Comments
Feige110                          Std              00:22      all
PSNJ01591443+1901339   Follow          01:00     i                          A
LSQ13ddu                        Follow          01:32     g16           1         A/B
OGLE-2013-SN-131          Class           02:51      g13           1         A/B        contamination by host, hard to extract
SN2013gr                         Follow          03:18     g11/g16     1         A            seeing 0.9"
LSQ13ddu                        Follow          05:06     g11           1         A            Ic or IIn?  Moon set 05:46
LSQ13dkp                        Class           06:17     g13           1         Z            WRONG OBJECT. Confusing finding chart.
LSQ13dkp                        Class           06:48     g13           1         A            Redone. young Ia at z~0.075. Seeing 0.85"
CATA50013A                    Class           07:02     g13           1         B            Probable IIP z~ 0.03
LSQ13deg                        Class           07:24     g13           1         A
LTT3864                           Std              08:15     all          all

Comments